Circular No. 8279

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 B1 (LINEAR) An apparently asteroidal object reported by the LINEAR project (dis-

covery observation below), and posted on the NEO Confirmation Page, has been reported by J. Young (from CCD exposures taken with the 0.6-m reflector at Table Mountain Observatory on Jan. 30.1 UT) to have a 3′′ round coma with a broad extension 4′′ long in p.a. 300^{o}–320^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Jan. 29.15539  5^{h}12^{m}54.^{s}24  +0^{o}59′16′′.2  19.1

The available astrometry (including prediscovery observations made on Jan. 28.4 by the NEAT project at Haleakala), the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-B73.

T = 2006 May 9.827 TT  ω = 322.^{o}421
Ω = 277.051 }2000.0
q = 2.15247 AU  i = 123.270

SATURN IX (PHOEBE) J. Bauer, D. Simonelli, and B. Buratti, Jet Propulsion Laboratory,

California Institute of Technology, report that photometric observations of Saturn IX, obtained with the Table Mountain Observatory 0.6-m telescope on 2003 Dec. 2–3 and 2004 Jan. 13–16 UT in support of NASA’s Cassini mission, show a magnitude variation of 0.12 in Johnson V and yield an updated period of 9.2735 ± 0.0015 hr (a factor-of-10 improvement over prior publications). The light-curve peak was at 2004 Jan. 14.308 UT, which they find corresponds to longitude 355^{o} ± 15^{o} (sub-observer point).

U SCORPII B. E. Schaefer, Louisiana State University, writes that a previously

unknown eruption of the recurrent nova U Sco was discovered on Harvard College Observatory archival photographs: plate AC18624 shows U Sco at B = 9.1 (near maximum brightness) on 1917 Mar. 6. With eruptions of U Sco having occurred in 1906, 1917, 1936, 1945, 1969, 1979, 1987, and 1999, Schaefer notes that U Sco has a fairly constant recurrence cycle of 8–12 yr — with ∼ 25 percent of the outbursts being missed due to proximity to the sun (including potential missed outbursts around 1926 and 1957) — adding that the next U Sco eruption should occur sometime during 2007–2011.

2004 January 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green